4) Draw the shear force and bending moment diagrams for the beam subjected to the loading shown. 30) KN SUKN
Answer #1

30 KN SOKN Reactions: Efy -0 => i 30 +50 : 80 KN RD - 126 KN EMB=0 . 50x3 = 30X1 + Ro X5 :: Ro = 24 KN from 0, RB + 24 = 80 KPoint of contra flexure (E):_ It is the point at which B.M.=0. EM=0 -30 x ze + 56 x(x-1)=0 -30 * +56 * - 56 -0 .: X: 2.15 m f